A simple pendulum osscilates with an average period of 1.25 sec . determine it's length
1
Expert's answer
2020-10-12T03:41:46-0400

A period of a simple pendulum is given by the following equation:


"T = 2\pi\sqrt{\dfrac{l}{g}}"

where "l" is the length of the pendulum and "g = 9.81m\/s^2" is the gravitational acceleration. Expressing "l", obtain:


"l = \dfrac{T^2g}{4\pi^2} = \dfrac{1.25^2\cdot 9.81}{4\pi^2} \approx 0.39m"

Answer. 0.39m.